OpenStudy (swaugh):

medal + fan What type of relationship is shown in the table? An x-y table shows the following coordinate pairs: two comma twenty-eight, three comma forty-two, and four comma fifty-six. quadratic inverse direct linear only

Directrix (directrix):

The y-values go up by 14 when x goes up by 1. That relationship is linear. 28 + 14 = 42; 42 + 14 = 56

Directrix (directrix):

As for " direct" I assume that this means direct variation. The form for an equation for direct variation is: y = k*x where k is the constant of variation. In this problem, k is the constant of variation.

Directrix (directrix):

@swaugh What do you think the complete answer is?